OpenStudy (anonymous):

simplify v^-5 . v^3 write answer with positive exponent only

OpenStudy (anonymous):

is there suppose to be a . in the middle

OpenStudy (anonymous):

yes i think it means multiply

OpenStudy (anonymous):

\[v^{-5}=\frac{ 1 }{ v^5 }\]

OpenStudy (anonymous):

v^−5v^3 v^−5+^3 1/^v2

OpenStudy (anonymous):

so\[v^{-5}\cdot v^3 = \frac{ 1 }{ v^5 }\cdot v^3 = \frac{ v^3 }{ v^5 }=\frac{ 1 }{ v^2 }\]

OpenStudy (anonymous):

so final answer is 1/v^2

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem:)